OpenStudy (anonymous):

The point P(18, 24) is on the terminal side of θ. Evaluate sin θ.

OpenStudy (anonymous):

help please :(

OpenStudy (triciaal):

welcome

OpenStudy (anonymous):

hi can u help me out?

OpenStudy (triciaal):

terminal means end tan of the angle y/x given (18, 24) tan = 24/18) =4/3 angles recorded from right horizontal anti-clockwise sin the ratio opposite/ hypotenuse quadrants with tan positive I and III

OpenStudy (anonymous):

oh so is the answer 4/3 or 4/5?

OpenStudy (triciaal):